OpenStudy (tia_sweetheart17):

solve for x. 7(x-7)=6(2+x)

OpenStudy (nincompoop):

solve the equation first by utilizing distributive property an example would be a(b + c) = 3(5a + b) --> (ab + ac) = (15a + 3b)

OpenStudy (daffy_duck):